write an equation of the line that passes through each pair of points. find the slope and y-intercept.
1. (1,1) (7,4)
2. (5,7) (0,6)
3. (-5/4,1) (-1/4,3/4)
how do you find the y - intercept ?

Answer by jake_6233(74)

1. (1,1) (7,4)
2. (5,7) (0,6)
3. (-5/4,1) (-1/4,3/4)
how do you find the y - intercept
We may only have room here for one answer if you send me an email I can work the other two y= mx+b is what you need to remember m= slope and b is the y intercept so (4-1)/(7-1) = 3/6 or 1/2 =m Then because each point in the problem is on the line substitue a pair of coordinates into the equation y -y1=m(x-x1)
y-4= 1/2(x-7) y= 1/2x - 7/2 + 8/2 y= 1/2x+ 1/2 (y=mx+b) b = 1/2
I hope that this helps